A baggage handler drops your 8.50 kg suitcase onto a conveyor belt running at 2.50 m/s . The materials are such that μs = 0.600 and μk = 0.290.

How far is your suitcase dragged before it is riding smoothly on the belt?
Answer #1

Acceleration, a = uk x g = 0.29 x 9.8 = 2.84 m/s^2

Using 2nd equation of motion,

V^2 = u^2 + 2as

2.5^2 = 0 + 2 x 2.84 x s

Distance, s = 1.10 m

Comment in case any doubt please rate my answer ....